# Guiding Time-Varying Generative Models with Natural Gradients on Exponential Family Manifold

### v1

@@ 5.2 Graphical Model Recovery with Informative Sufficient Statistics
In the covariate shift setting, we observe joint samples from the source $p_{X,Y}$, but only have target domain covariates from $q_{X}$. The goal is to find $\psi^{-1}$. We find the reverse process by minimizing $\mathrm{KL}[p_{X},q_{t}]$ using Algorithm 2, where $X_{0}\sim q_{0}$ are set to be the target domain covariates.
@@
We further test our algorithm on the Office+Caltech classification dataset [Gong et al., 2012] which is an object recognition dataset with photos collected from four different places: amazon, dslr, webcam, caltech. The task is to train a source classifier using one of the places, and test it on samples from another place. We test the performance of ntKiNG against two other particle-based transport methods WGF and MMD flow that also matches $q_{t}$ with $p_{X}$. The performance is measured by the percentage gains compared with directly applying the source classifier to the target samples. The results in Table 1 show that our method achieves the most accuracy gains comparing to WGF and MMD flow. In 10 out of 12 domain adaptations settings, ntKiNG improves the testing accuracy.
@@ 6 Related Works
Appendix A Infinite-dimensional Exponential Family
One important class of exponential family is infinite dimensional exponential family.
Example 4.
An infinite-dimensional exponential family is an exponential family with a sufficient statistic $T({\boldsymbol{x}}):=k({\boldsymbol{x}},\cdot)$:
$q({\boldsymbol{x}};\theta)=\exp\big{(}\langle\theta,k({\boldsymbol{x}},\cdot)\rangle-A(\theta)\big{)},$
where $A(\theta):=\int\exp\big{(}\langle\theta,k({\boldsymbol{x}},\cdot)\rangle-A(\theta)\big{)}\mathrm{d}{\boldsymbol{x}}$ and $\theta$ is in an RKHS $\mathcal{H}$ with kernel $k$.
Given a dataset of $n$ data points $\{{\boldsymbol{x}}_{i}\}_{i=1}^{n}\sim q_{\theta}$, the empirical natural gradient is defined as
Definition A.1.
$\mathrm{NGD}:=\mathop{\rm argmax}\limits_{f\in\mathcal{H}}\langle f,\nabla\mathcal{L}(\theta)\rangle_{\mathcal{H}}-\frac{1}{2}\langle f,\hat{{\boldsymbol{\Sigma}}}f\rangle_{\mathcal{H}}-\lambda\|f\|_{\mathcal{H}}^{2},$ (12)
where $\hat{{\boldsymbol{\Sigma}}}$ is the empirical covariance operator defined as
$\hat{{\boldsymbol{\Sigma}}}:=\frac{1}{n}\sum_{i=1}^{n}\phi({\boldsymbol{x}}_{i})^{\top}{\boldsymbol{H}}\phi({\boldsymbol{x}}_{i}),$
and ${\boldsymbol{H}}$ is a centering matrix and $\lambda>0$.
Due to the Representer Theorem, we can see that the optimal solution takes the form $\mathrm{NGD}:=\sum_{i}\alpha_{i}\phi({\boldsymbol{x}}_{i},\cdot)$.
Since our aim is to match the projection $\delta\in\mathcal{H}$ to the natural gradient, we let $\delta=\sum_{i=1}^{n}\alpha_{i}k({\boldsymbol{x}}_{i},\cdot)$. The regularized objective of Equation 4 takes the form
$\small J({\boldsymbol{\alpha}}):=$ $\int\lambda_{t_{0}}(t)\mathbb{E}\left[\left(s_{t}(X_{t})-\langle{\boldsymbol{\alpha}},{\boldsymbol{k}}(X_{t})-\mathbb{E}[{\boldsymbol{k}}(X_{t}^{\prime})]\rangle\right)^{2}\right]\mathrm{d}t+\lambda{\boldsymbol{\alpha}}{\boldsymbol{K}}{\boldsymbol{\alpha}}$ (13)
where ${\boldsymbol{k}}(X):=[k({\boldsymbol{x}}_{1},X),k({\boldsymbol{x}}_{2},X),\dots]$. The limiting solution of Equation 13 is
$\lim_{\sigma\to 0}{\boldsymbol{\alpha}}_{t_{0}}({\boldsymbol{w}})=\left(\mathrm{Cov}[{\boldsymbol{k}}(X_{t_{0}})]+\lambda{\boldsymbol{K}}\right)^{-1}\mathbb{E}\left[\nabla{\boldsymbol{k}}(X_{t_{0}}){\boldsymbol{h}}^{\top}(X_{t_{0}};{\boldsymbol{w}})\right].$
Substituting the form of $\mathrm{NGD}$ and solving (12) for ${\boldsymbol{\alpha}}$, we get
${\boldsymbol{\alpha}}^{*}=\left(\mathrm{Cov}[{\boldsymbol{k}}(X_{t_{0}})+\lambda{\boldsymbol{K}}]\right)^{-1}\left(\mathbb{E}_{p}[{\boldsymbol{k}}({\boldsymbol{x}})]-\mathbb{E}_{q}[{\boldsymbol{k}}({\boldsymbol{x}})]\right).$
Thus, the optimal drift parameter could be obtained via ${\boldsymbol{w}}^{*}:=\mathop{\rm argmin}\limits_{\boldsymbol{w}}\|{\boldsymbol{\alpha}}^{*}-\lim_{\sigma\to 0}{\boldsymbol{\alpha}}_{t_{0}}({\boldsymbol{w}})\|^{2}$, which has a similar solution to Equation 9 or Equation 11 depending on which model is used for the drift.
Appendix B Proof of Theorem 3.1
Theorem B.1.
@@ Appendix A Infinite-dimensional Exponential Family
Notice that Equation 14 is a quadratic minimization problem. Let $B_{t}=T(X_{t})-\mathbb{E}[T(X^{\prime}_{t})]$. Firstly, we find the derivative of the quadratic term. For each $t$, since $\nabla_{\boldsymbol{\delta}}\,\bigl{(}\langle{\boldsymbol{\delta}},B_{t}\rangle^{2}\bigr{)}=2\,\langle{\boldsymbol{\delta}},B_{t}\rangle\,B_{t},$ it follows that
@@
To find the minimizer, set this gradient to zero:
@@
Appendix C Proof of Theorem 4.2
@@
Thus, due to Lemma C.1, we have
@@
Appendix D Proof of Theorem 4.3
@@
Recall, that we try to minimize Equation 8
$\|\nabla\mathcal{L}({\boldsymbol{\theta}}_{t_{0}})-\mathbb{E}[\nabla T(X_{t_{0}})\nabla{\boldsymbol{h}}_{w}(X_{t_{0}})]\|_{\mathcal{F}_{t_{0}}^{-1}}^{2}+\lambda\|w\|^{2}_{\mathcal{H}},$ (18)
@@
Notice the product $\mathbb{E}[\nabla T(X_{t_{0}})\nabla k(X_{t_{0}},\cdot)]\mathbb{E}[\nabla T(X_{t_{0}})\nabla k(X_{t_{0}},\cdot)]^{\top}=\mathbb{E}[\nabla T(X_{t_{0}})\nabla\nabla k(X_{t_{0}},X_{t_{0}}^{\prime})\nabla^{\top}T(X_{t_{0}}^{\prime})]$, we obtain the desired result.
Appendix E Stein Exponential Family
@@
This parametric model opens up applications such as variational inference and sampling: Given an unnormalized density $p$, we want to train a generative model to sample from $p$. We can design a Stein exponential family so that we can approximately minimize $\mathrm{KL}[p,q_{t}]$ using kernel NGD by pushing particles towards the target distribution.
Appendix F Reverse KL Wasserstein Gradient Flow and MMD Flow
The WGF dynamics that minimize $\mathrm{KL}[q_{t},p]$ move particles $X_{t}$ using a simple velocity field:
@@ Appendix F Reverse KL Wasserstein Gradient Flow and MMD Flow
where the gradient of log density could be easily estimated via kernel density estimation and the MMD flow minimizes $\mathrm{MMD}[X_{1},X_{t}]$ using the following velocity field:
@@
Appendix G Experiment Setup in Section 5
@@
G.1 Comparison with Reverse KL Wasserstein Gradient Flow and MMD Flow
G.1.1 Dataset and Pre-processing
@@ G.1.2 Parameter Tuning
G.2 Graphical Model Recovery
G.2.1 Dataset and Pre-processing
@@ G.2.2 Parameter Tuning
G.3 Experiment 2: Covariate Shift
G.3.1 Dataset and Pre-processing

### v2

@@ 5.2 Graphical Model Recovery with Informative Sufficient Statistics
In the covariate shift setting, we observe joint samples from the source $p_{X,Y}$, but only have target domain covariates from $q_{X}$. The goal is to find $\psi^{-1}$. We find the reverse process by minimising $\mathrm{KL}[p_{X},q_{t}]$ using Algorithm 2, where $X_{0}\sim q_{0}$ are set to be the target domain covariates.
@@
We further test our algorithm on the Office+Caltech classification dataset [Gong et al., 2012] which is an object recognition dataset with photos collected from four different places: amazon, dslr, webcam, caltech. The task is to train a source classifier using one of the places, and test it on samples from another place. We reduce the dimension of the dataset to 50 using PCA and test the performance of ntKiNG against two other particle-based transport methods WGF and MMD flow that also matches $q_{t}$ with $p_{X}$. The performance is measured by the percentage gains compared with directly applying the source classifier to the target samples. The results in Table 1 show that our method achieves the most accuracy gains comparing to WGF and MMD flow. In 10 out of 12 domain adaptations settings, ntKiNG improves the testing accuracy.
@@
5.4 Sample Denoising with Pretrained Energy-based Model (EBM)
Figure 8: Denoised samples obtained by ntKiNG and WGF. ntKiNG uses pretrained energy-based model as $T$.
Figure 9: Denoised MNIST images
To further investigate the performance of iNGD on higher-dimensional datasets and more sophisticated sufficient statistics, we conduct two denoising experiments.
We begin by drawing samples $Y\sim q$ and train a deep energy-based model using denoising score matching [Song and Kingma, 2021]. We define the sufficient statistic $T({\boldsymbol{x}})$ as the output of the penultimate layer of the EBM, ensuring that $T({\boldsymbol{x}})$ captures information about the distribution $q$. The structure of the EBM can be found in Section F.4.
We then construct noisy samples by adding Gaussian noises to $Y$ as $X_{0}$ and run ntKiNG and WGF using $Y$ and $X_{0}$ as the target and initial samples with 1000 iterations.
We first test the procedure on an S-curve dataset and the result is shown in Figure 8. It can be seen that ntKiNG more effectively recovers the S-shape compared to WGF.
We then perform the experiment on the MNIST dataset (downsampled to 18 by 18), using the same model architecture and estimator to construct $T$. As we can see from Figure 9, both methods act as decent denoisers, but ntKiNG produces cleaner reconstructions than WGF.
@@ 6 Related Works
7 Limitations and Future Works
The effectiveness of the guidance heavily relies on the choice of the exponential family manifold, which is determined by the sufficient statistics $T$. If the guidance is weak, the particles may fail to converge to the target distribution, as demonstrated in Figure 3. In this paper, we show that sophisticated sufficient statistics—such as RBF features or a pretrained EBM—can achieve promising results. Developing theories to better understand the choices of sufficient statistics is an important future work.
In this paper, we only focus on drift-based generative model for its simplicity. An interesting future work is studying the applicability of our methods to other types of generative models (e.g., GAN or diffusion model).
The primary computational bottleneck of our method lies in the inversion of the $\mathrm{dim}(T)\times\mathrm{dim}(T)$ matrix ${\boldsymbol{\Gamma}}$ in (11), which become computationally prohibitive if $\mathrm{dim}(T)$ is high. For some choices of $T$ (e.g., RBF), reducing $\mathrm{dim}(T)$ also reduces its expressiveness. Thus, extending iNGD to a high-dimensional $T$ is an urgent future work. Both our method and MMD flow requires computing kernel matrix, which has an $n^{2}$ computational complexity. However, there is no matrix inversion involved in MMD flow.
More generally, despite the promising results, the benefits of guiding the generative model in a parametric space remain to be clarified. Developing theories and applications that compare our method with established generative approaches—such as diffusion models—represents an interesting direction for future work.
Acknowledgements
We thank four anonymous reviewers and the area chair for their insightful comments. We thank Dr. Sam Power, Dr. Katerina Karoni and Bristol Machine Learning Reading Group for helpful discussions.
Appendix A Proof of Theorem 3.1
Theorem A.1.
@@ Acknowledgements
Notice that Equation 12 is a quadratic minimisation problem. Let $B_{t}=T(X_{t})-\mathbb{E}[T(X^{\prime}_{t})]$. Firstly, we find the derivative of the quadratic term. For each $t$, since $\nabla_{\boldsymbol{\delta}}\,\bigl{(}\langle{\boldsymbol{\delta}},B_{t}\rangle^{2}\bigr{)}=2\,\langle{\boldsymbol{\delta}},B_{t}\rangle\,B_{t},$ it follows that
@@
To find the minimiser, set this gradient to zero:
@@
Appendix B Proof of Theorem 4.2
@@
Thus, due to Lemma B.1, we have
@@
Appendix C Proof of Theorem 4.3
@@
Recall, that we try to minimise Equation 8
$\|\nabla\mathcal{L}({\boldsymbol{\theta}}_{t_{0}})-\mathbb{E}[\nabla T(X_{t_{0}})\nabla{\boldsymbol{h}}_{w}(X_{t_{0}})]\|_{\mathcal{F}_{t_{0}}^{-1}}^{2}+\lambda\|w\|^{2}_{\mathcal{H}}.$ (16)
Expanding the first square, up to a constant that does not depend on $w$, we obtain
$\mathbb{E}[\nabla T(X_{t_{0}}){\boldsymbol{h}}_{w}(X_{t_{0}})]^{\top}\mathcal{F}_{t_{0}}^{-1}\mathbb{E}[\nabla T(X_{t_{0}}){\boldsymbol{h}}_{w}(X_{t_{0}})]-2\nabla^{\top}\mathcal{L}({\boldsymbol{\theta}}_{t_{0}})\mathcal{F}_{t_{0}}^{-1}\mathbb{E}[T(X_{t_{0}}){\boldsymbol{h}}_{w}(X_{t_{0}})]+\lambda\|w\|^{2}_{\mathcal{H}}.$ (17)
By definition, ${\boldsymbol{h}}_{w}(X_{t_{0}})=\langle w,\nabla k(X_{t_{0}},\cdot)\rangle$, we obtain a quadratic form with respect to $w$,
$\langle w,\mathbb{E}[\nabla T(X_{t_{0}})\nabla k(X_{t_{0}},\cdot)]^{\top}\mathcal{F}_{t_{0}}^{-1}\mathbb{E}[\nabla T(X_{t_{0}})\nabla k(X_{t_{0}},\cdot)]w\rangle-\langle w,2\mathbb{E}[T(X_{t_{0}})\nabla k(X_{t_{0}},\cdot)]^{\top}\mathcal{F}_{t_{0}}^{-1}\nabla\mathcal{L}({\boldsymbol{\theta}}_{t_{0}})\rangle+\lambda\|w\|^{2}_{\mathcal{H}},$ (18)
where we used $a^{\top}BCd=(BCd)^{\top}a=d^{\top}C^{\top}B^{\top}a=\langle d,C^{\top}B^{\top}a\rangle$ on the second term and the fact that the inverse of Fisher Information Matrix $\mathcal{F}_{t_{0}}$ is a symmetric matrix. Differentiating both sides by $w$ and setting the gradient to zero, we obtain the following optimality condition of the least squares:
$0\displaystyle 2\mathbb{E}[\nabla T(X_{t_{0}})\nabla k(X_{t_{0}},\cdot)]^{\top}\mathcal{F}_{t_{0}}^{-1}\mathbb{E}[\nabla T(X_{t_{0}})\nabla k(X_{t_{0}},\cdot)]w-2\mathbb{E}[T(X_{t_{0}})\nabla k(X_{t_{0}},\cdot)]^{\top}\mathcal{F}_{t_{0}}^{-1}\nabla\mathcal{L}({\boldsymbol{\theta}}_{t_{0}})+2\lambda w=$ 0 .
@@
Notice the product $\mathbb{E}[\nabla T(X_{t_{0}})\nabla k(X_{t_{0}},\cdot)]\mathbb{E}[\nabla T(X_{t_{0}})\nabla k(X_{t_{0}},\cdot)]^{\top}=\mathbb{E}[\nabla T(X_{t_{0}})\nabla\nabla k(X_{t_{0}},X_{t_{0}}^{\prime})\nabla^{\top}T(X_{t_{0}}^{\prime})]$, where $\nabla\nabla k(x,y):=\nabla_{x}\nabla_{y}k(x,y)$ we obtain the desired result.
Appendix D Stein Exponential Family
@@
This parametric model opens up applications such as variational inference and sampling: Given an unnormalized density $p$, we want to train a generative model to sample from $p$. We can design a Stein exponential family so that we can approximately minimise $\mathrm{KL}[p,q_{t}]$ using kernel NGD by pushing particles towards the target distribution.
Appendix E Reverse KL Wasserstein Gradient Flow and MMD Flow
The WGF dynamics that minimise $\mathrm{KL}[q_{t},p]$ move particles $X_{t}$ using a simple velocity field:
@@ Appendix E Reverse KL Wasserstein Gradient Flow and MMD Flow
where the gradient of log density could be easily estimated via kernel density estimation and the MMD flow minimises $\mathrm{MMD}[Y,X_{t}]$ using the following velocity field:
@@
Appendix F Experiment Setup in Section 5
@@
F.1 Comparison with Reverse KL Wasserstein Gradient Flow and MMD Flow
F.1.1 Dataset and Pre-processing
@@ F.1.2 Parameter Tuning
F.2 Graphical Model Recovery
F.2.1 Dataset and Pre-processing
@@ F.2.2 Parameter Tuning
F.3 Experiment 2: Covariate Shift
F.3.1 Dataset and Pre-processing
@@ F.3.2 Parameter Tuning
F.4 Experiment 3: Denoising
The deep EBM is constructed using the following pyTorch code:
# Define the MLP-based energy-based model
class EnergyBasedModel(nn.Module):
    def __init__(self, input_dim):
        super(EnergyBasedModel, self).__init__()
        self.network = nn.Sequential(
            nn.Linear(input_dim, 1024),
            nn.SiLU(),
            nn.Linear(1024, 1024),
            nn.SiLU(),
            nn.Linear(1024, 211),
            nn.SiLU(),
            nn.Linear(211, 1)
    def forward(self, x, penultimate=False, flattened=False):
        x = x.view(x.size(0), -1)  # Flatten the input
        if not penultimate:
            return self.network(x)
        else:
            for i, layer in enumerate(self.network):
                x = layer(x)
                if i == len(self.network) - 2:
                    break
            return x
The model is trained using 10000 samples of $Y$, with a batch size 777, and adam optimizer with a step size of 0.001. In the S-curve experiment, we add a Gaussian noise to the target sample $Y$ with a standard deviation 0.3, and in the MNIST experiment, we add a Gaussian noise to $Y$ with a standard deviation 0.2.
The noise used in denoising score matching mathces the noise added to $Y$.